# Eulerian magnitude homology: diagonality, injective words, and regular path homology

### v1

@@ Introduction
Let $G$ be a finite regularly diagonal undirected graph with diagonal eulerian magnitude homology in all bidegrees $(k,k)$ with $0\leq k\leq n$. Then, $G$ is the complete graph on $n+1$ vertices.
As all complete undirected graphs are regularly diagonal, this shows that eulerian magnitude homology detects the complete graphs. We recall here that this is not the case for magnitude homology, for which the family of undirected diagonal graphs is quite large. When we consider directed graphs the story is different, and we do not know of any characterization of regularly diagonal digraphs. On the other hand, in analogy with what happens for magnitude homology of graphs, cones and, more generally, joins of regularly diagonal digraphs are regularly diagonal (see Theorem 3.9). We shall show in Section 3 that trees, bipartite graphs with alternating orientations, and transitive tournaments, are other examples of regularly diagonal digraphs.
Besides the combinatorial appeal, our interest in (regular) diagonality of graphs stems from the fact that it is extremely useful in computations of regular path homology [HY15, TY22], a variant of the more classical and established path homology [GLMY20]. With an eye towards these computations, we develop the analogue of the magnitude-path spectral sequence [Asa23] – which we call the regular magnitude-path spectral sequence. Using the indices also used in [Asa23], we have:
@@
    The spectral sequence converges to the homology of the complex of injective words on $G$.
@@
The regular magnitude-path spectral sequence is the spectral sequence associated to the filtered simplicial set $N^{\iota}(\mathbf{Reach}_{G})$, which is the injective nerve of the reachability category [CR24] of a directed graph $G$. The injective nerve is defined as the simplicial set on functors to $\mathbf{Reach}_{G}$ that are injective on objects – see Section 5.1 – and it agrees with the directed flag complex of the transitive closure $\mathrm{Pre}(G)$ (also known as reachability preorder [DIMZ24, HR23]) of $G$. In particular, $N^{\iota}(\mathbf{Reach}_{G})$ models the complex of injective words on $G$ – see Definition 4.3 – which in turn generalizes the classical complex of injective words [Far79, BW83, CLM20]. This leads to an interpretation of eulerian magnitude homology in terms of directed flag complexes and reachability preorders, which we believe to be of independent interest. As applications, we show the following:
@@
Let $G$ be a finite and connected regularly diagonal directed graph. Then, the regular path homology of $G$ is the homology of the directed flag complex $\mathrm{dFl}(\mathrm{Pre}(G))$.
As a consequence, the reduced regular path homology of trees and transitive tournaments is trivial, whereas the regular path homology of complete graphs is the homology of a wedge of spheres. This last computation allows us to prove that regular path homology (and more generally, the pages of the regular magnitude-path spectral sequence) are not homotopy invariants – see Corollary 5.18.
@@
The paper is organized as follows. We recall in Section 1 some general background about first standard and eulerian magnitude homology, and further standard and regular path homology. We proceed by investigating in Section 2 diagonality conditions for eulerian magnitude homology of undirected graphs, and we then expand this study in Section 3 examining diagonality properties of eulerian magnitude homology of digraphs.
In Section 4 we extend to the setting directed graphs the notion of complex of injective words, which will emerge as the target object of the regular magnitude-path spectral sequence developed in Section 5.
@@
All (directed) graphs are finite. Undirected graphs are simple. In the case of directed graphs, we shall not allow multiple directed edges, which means that the set of edges of a directed graph $G$ with vertices $V(G)$ is a subset of the ordered pairs $V(G)\times V(G)$. All modules are taken over a base commutative ring $R$. We shall use the following notations:
$\rho(G)$ the directed graph associated to an undirected graph $G$ cf. Notation 1.1
@@ 1. Eulerian magnitude homology and regular path homology
be the free $R$-module on the paths of length $l$ consisting of $k+1$ vertices of $G$. The differential $\partial=\partial_{k,l}\colon\mathrm{MC}_{k,l}(G;R)\to\mathrm{MC}_{k-1,l}(G;R)$ acts on $(k+1)$-uples $(x_{0},\dots,x_{k})$ of $G$ by
@@
forgets the $i$-th entry if the length is preserved, and it is set to be 00 otherwise. The pair $(\mathrm{MC}_{*,l}(G;R),\partial)$ is a chain complex by [HW17, Lemma 2.11]. The magnitude homology $\mathrm{MH}_{*,*}(G;R)$ of $G$ is the bigraded $R$-module $\bigoplus_{k,l\geq 0}\mathrm{MH}_{k,l}(G;R)$, where
$\mathrm{MH}_{k,l}(G;R)\coloneqq H_{k}(\mathrm{MC}_{*,l}(G;R),\partial)$
@@
The definition of magnitude homology extends verbatim to directed graphs – cf[Asa23].
@@
Let $G$ be a graph on the set of vertices $V$. The $(k,\ell)$-eulerian magnitude chain $\mathrm{EMC}_{k,\ell}(G)$ is the free $R$-module generated by tuples $(x_{0},\dots,x_{k})\in V^{k+1}$ of length $\ell$ such that $x_{i}\neq x_{j}$ for every $0\leq i,j\leq k$.
@@
Let $G$ be a (directed) graph with vertex set $V$. Then, for each positive integer $\ell\geq 0$, we have
$\mathrm{EMC}_{\ast,\ell}(G)=\bigoplus_{a,b\in V}\mathrm{EMC}_{\ast,\ell}(a,b)\ ,$
@@
Eulerian magnitude homology does not satisfy Künneth. In fact, let $I_{2}$ be the linear undirected graph with two vertices $x_{0},x_{1}$ and an undirected edge $\{x_{0},x_{1}\}$. Its eulerian magnitude homology is non-trivial only in bidegrees $(k,l)=(0,0),(1,1)$, in which cases it is of rank 2. The Cartesian product $I_{2}\times I_{2}$ with itself is the cycle graph $C_{4}$ on four vertices, which has non-trivial eulerian magnitude homology in bidegree $(k,l)=(3,5)$. By dimensionality, the generators in bidegree $(3,5)$ can not come from generators of $\mathrm{EMH}_{*,*}(I_{2})\otimes\mathrm{EMH}_{*,*}(I_{2})$.
@@
By Proposition 1.8, eulerian magnitude homology contains torsion. On the other hand, such torsion is “bounded” if we restrict to (the category of) graphs with bounded genus, as we shall explain. To state the result, we need a cohomological version of eulerian magnitude homology. Recall that for a graph $G$, its eulerian magnitude cochain complex is defined as
$\mathrm{EMC}^{k}_{\ell}(G;\mathbb{Z})\coloneqq\mathrm{Hom}(\mathrm{EMC}_{k,\ell}((G);\mathbb{Z}),\mathbb{Z})\ ,$
equipped with the dual differential. The eulerian magnitude cohomology of $G$ is then the cohomology of the cochain complex $\mathrm{EMC}^{*}_{\ell}(G;\mathbb{Z})$. Eulerian magnitude homology and eulerian magnitude cohomology are related by a short exact sequence from the universal coefficients theorem, similar to the short exact sequence relating magnitude homology and magnitude cohomology [Hep22, Remark 2.5].
@@
For every pair of integers $k,g\geq 0$, there exists $m=m(g,k)\in\mathbb{Z}$ which annihilates the torsion subgroup of $\mathrm{EMH}^{k}_{*}(G;\mathbb{Z})$, for each (directed) graph $G$ of genus at most $g$.
@@
The analogous statements of [CC24, Corollary 2.12] and [CCR24, Theorem 5.4] can be adapted verbatim to the case of eulerian magnitude cohomology of both directed and undirected graphs; showing that eulerian magnitude cohomology, as a functor on the category of (directed) graphs with contractions, is finitely generated. Then, the result follows with the same proof of [CC24, Theorem 3.5] (see also [MR23, Theorem 3.20]). ∎
Proposition 1.9 holds because eulerian magnitude cohomology, seen as a functor, is a finitely generated module over the category of directed graphs (or, more generally, quivers). Another consequence of this fact is the following bound on the rank growth:
Proposition 1.10.
Let $\mathbb{K}$ be a field, and $g\geq 0$. Then, there exists a polynomial $f\in\mathbb{Z}[t]$ of degree at most $g+k+1$, such that, for all graphs $G$ of genus at most $g$, we have
$\dim_{\mathbb{K}}\mathrm{EMH}_{*}^{k}(G;\mathbb{K})\leq f(|E(G)|)\ ,$
where $|E(G)|$ is the number of edges of $G$.
Proof.
The proof follows the arguments of [CC24, Corollary 3.1]. ∎
We provide some examples of computation.
@@
Let $I_{n}$ be the linear undirected graph on $n$ vertices. Then, the maximum index $\ell$ for which $\mathrm{EMH}_{k,\ell}(I_{n})\neq 0$ is $\ell=\frac{n(n-1)}{2}$, achieved for $k=n-1$.
@@
Let $G=I^{n}_{2}$ be the $n$-fold Cartesian product of the graph $I_{2}$ with itself. The maximum index $\ell$ for which $\mathrm{EMH}_{k,\ell}(I^{n}_{2})\neq 0$ is the maximum total Hamming distance between pairs of consecutive elements in any permutation of all $2^{n}$ binary words of length $n$ – see the A271771 sequence.
@@
$\ell(n)=\begin{cases}\frac{n^{2}}{2}-n+1&\text{ if }n\text{ is even}\\ \frac{(n-1)^{2}}{2}&\text{ if }n\text{ is odd}\end{cases}$
and is achieved in bidegree $(n-1,\ell(n))$. Furthermore, for every $(k,\ell)$ it holds that $\mathrm{EMH}_{k,\ell}(C_{n})$ is upper bounded by
$2\cdot\binom{n}{k+1}\binom{\ell-1}{k}\ ,$
@@
We consider a directed graph $G$ on the vertex set $V$. For $n\geq 0$, an elementary path is an ordered sequence $(x_{0},\dots,x_{n})$ of vertices of $G$. We let $\Lambda_{n}(G)$ be the module freely generated by the $n$-elementary paths of $G$. There is a differential $\partial\colon\Lambda_{n}(G)\to\Lambda_{n-1}(G)$ as the alternating sums of the face maps forgetting the $i$-th entry of the elementary paths. An elementary path $(x_{0},\dots,x_{n})$ is called allowed if $(x_{i},x_{i+1})$ is a directed edge of $G$ for all $i=0,\dots,n-1$.
@@ 2. Regularly diagonal graphs
In this section we focus on the diagonality properties of eulerian magnitude homology of graphs, providing a complete characterization of diagonal undirected graphs for eulerian magnitude homology.
@@
sending a tuple of $G$ to the tuple in $G^{\mathrm{op}}$ with same vertices and opposite order. Then, this map induces an isomorphism of magnitude homology groups [IM24, Lemma 6.3], and, likewise, of eulerian magnitude homology groups. ∎
@@
If $G$ has no directed cycles, then the eulerian magnitude chain complex of $G$ is the magnitude chain complex of $G$. ∎
Lemma 2.6.
@@
We shall see that Lemma 2.7 can be improved, and in Corollary 2.13 we provide a complete characterization of regularly diagonal undirected graphs. To proceed with such characterization, we first need to analyze the relation between the existence of cycles in an undirected graph and its eulerian magnitude homology. First, we recall the following fact – see [GM24, Lemma 3.2]:
@@
Recall from Example 2.1 that undirected trees are diagonal graphs, and, as trees have no cycles, the eulerian magnitude homology in bidegree $(k,k)$ $k\geq 2$ is trivial by Proposition 2.9. One may be tempted to infer that trees are also regularly diagonal; however, this is not true, as the next examples show:
@@
Let $G$ be a finite regularly diagonal undirected graph with diagonal eulerian magnitude homology in all bidegrees $(k,k)$ with $0\leq k\leq n$. Then, $G$ is the complete graph on $n+1$ vertices.
@@
  • pick any vertex $x_{0}$ of $G$ with $\ell(x)=D$;
@@
The defined tuple $(x_{0},\dots,x_{n})$ maximizes the length of all possible tuples of all vertices of $G$. In particular, its length $l$ is greater than $n$. As $(x_{0},\dots,x_{n})$ is also of maximal size, it is not in the image of the differential. We want to show that it is a cycle, concluding the proof. To show that the tuple $(x_{0},\dots,x_{n})$ is a cycle, it is enough to show that all subtuples $(x_{0},\dots,\hat{x_{i}},\dots,x_{n})$ have length smaller than $(x_{0},\dots,x_{n})$. For a series of vertices $x_{i-1},x_{i},x_{i+1}$ in the tuple, observe that we always have $d_{G}(x_{i-1},x_{i+1})\leq d_{G}(x_{i-1},x_{i})+d_{G}(x_{i},x_{i+1})$ by the triangle inequality. In addition, $d_{G}(x_{i-1},x_{i+1})\leq\ell(x_{i-1})$ by definition of $\ell(x_{i-1})$, and, by construction of the tuple $(x_{0},\dots,x_{n})$, we have that $d_{G}(x_{i-1},x_{i})$ is maximal among the distances $d_{G}(x_{i-1},y)$ with $y\in V(G)\setminus\{x_{0},\dots,x_{i-1}\}$. Hence, $d_{G}(x_{i-1},x_{i})\geq d_{G}(x_{i-1},x_{i+1})$. As a consequence, we have $d_{G}(x_{i-1},x_{i+1})<d_{G}(x_{i-1},x_{i})+d_{G}(x_{i},x_{i+1})$; that is, $\mathrm{len}(x_{0},\dots,\hat{x_{i}},\dots,x_{n})<\mathrm{len}(x_{0},\dots,x_{n})$ and $(x_{0},\dots,x_{n})$ is a cycle in $\mathrm{EMH}_{n,l}(G)$. However, this is not possible because $l>n$ and $G$ is regularly diagonal. ∎
@@
If $\mathrm{gir}(G)=\gamma$ then $G$ contains $C_{\gamma}$ as a subgraph. Then by Example 1.15 the statement follows. ∎

### v2

@@ Introduction
Let $G$ be a finite connected regularly diagonal undirected graph, with eulerian magnitude homology groups $\mathrm{EMH}_{k,k}(G)=0$ in all bidegrees $(k,k)$ for $k>n$. Then, $G$ is the complete graph on $n+1$ vertices.
As all complete undirected graphs are regularly diagonal, this shows that eulerian magnitude homology detects the complete graphs. We recall here that this is not the case for magnitude homology, for which the family of undirected diagonal graphs is quite larger. When considering directed graphs, the story is different, and we do not know of any characterization of regularly diagonal digraphs. On the other hand, in analogy with what happens for magnitude homology of graphs, cones and, more generally, joins of regularly diagonal digraphs are regularly diagonal (see Theorem 3.9). We shall show in Section 3 that directed trees, bipartite graphs with alternating orientations, and transitive tournaments, are other examples of regularly diagonal digraphs.
Besides the combinatorial appeal, our interest in (regular) diagonality of graphs stems from the fact that it is useful in computations of regular path homology [HY15, TY22], a variant of the more classical and established path homology [GLMY20]. With an eye towards these computations, we develop the analogue of the magnitude-path spectral sequence [Asa23] – which we call the regular magnitude-path spectral sequence. Using the same notation used in [Asa23], we have:
@@
    It converges to the homology of the complex of injective words on $G$.
@@
The regular magnitude-path spectral sequence is the spectral sequence associated to the filtered simplicial set $N^{\iota}(\mathbf{Reach}_{G})$, which is the injective nerve of the reachability category of a directed graph $G$ [CR24]. The injective nerve is defined as the simplicial set on functors to $\mathbf{Reach}_{G}$ that are injective on objects – see Section 5.1 – and it agrees with the directed flag complex of the transitive closure $\mathrm{Pre}(G)$ (also known as reachability preorder [DIMZ24, HR23]) of $G$. In particular, $N^{\iota}(\mathbf{Reach}_{G})$ models the complex of injective words on $G$ – see Definition 4.3 – which in turn generalizes the classical complex of injective words [Far79, BW83, CLM20]. This leads to an interpretation of eulerian magnitude homology in terms of directed flag complexes and reachability preorders, and, in turn, to a description of the directed flag complex in terms of injective nerves, which we believe to be of independent interest. As applications, we show the following:
@@
Let $G$ be a finite and connected regularly diagonal directed graph. Then, the regular path homology of $G$ is isomorphic to the homology of the complex of injective words on $G$.
As a consequence, the reduced regular path homology of directed trees and transitive tournaments is trivial, whereas the regular path homology of complete graphs is the homology of a wedge of spheres. This last computation allows us to prove that regular path homology (and more generally, the pages of the regular magnitude-path spectral sequence) are not homotopy invariants – see Corollary 5.20.
@@
The characterization of complete undirected graphs as by Theorem 5.15 was more recently shown also in [MIS25] by Sazdanovic and Martin. The similarities with their work arose from independent work.
@@
The paper is organized as follows. We recall in Section 1 some general background about first standard and eulerian magnitude homology, and further standard and regular path homology. We proceed by investigating in Section 2 diagonality conditions for eulerian magnitude homology of undirected graphs, and we then expand this study in Section 3 examining diagonality properties of eulerian magnitude homology of digraphs. In Section 4 we extend to the setting directed graphs the notion of complex of injective words, which will emerge as the target object of the regular magnitude-path spectral sequence developed in Section 5.
@@
Unless otherwise specified, all (directed) graphs and categories are finite. Undirected graphs are simple. In the case of directed graphs, we shall not allow multiple directed edges, which means that the set of edges of a directed graph $G$ with vertices $V(G)$ is a subset of the ordered pairs $V(G)\times V(G)$. All modules are taken over a base commutative ring $R$. We shall use the following notations:
$\rho(G)$ the directed graph associated to an undirected graph $G$; cf. Notation 1.1
@@
$\gamma_{n}^{s}$ girth of the maximal sub-cycle achievable by removing $s$ edges from $K_{n}$;
cf. Notation 2.21.
@@ Acknowledgments
The authors are thankful to Radmila Sazdanovic and Patrick Martin for a useful discussion that led to the proof of Proposition 1.8. The authors are also grateful to the anonymous referee, whose comments led to an improvement of the paper and to Remark 5.21.
@@ 1. Eulerian magnitude homology and regular path homology
be the free $R$-module on the tuples of length $l$ consisting of $k+1$ vertices of $G$. The differential $\partial=\partial_{k,l}\colon\mathrm{MC}_{k,l}(G;R)\to\mathrm{MC}_{k-1,l}(G;R)$ acts on $(k+1)$-uples $(x_{0},\dots,x_{k})$ of $G$ by
@@
forgets the $i$-th entry if the length is preserved, and it is set to be 0 otherwise. The pair $(\mathrm{MC}_{*,l}(G;R),\partial)$ is a chain complex by [HW17, Lemma 2.11]. The magnitude homology $\mathrm{MH}_{*,*}(G;R)$ of $G$ is the bigraded $R$-module $\bigoplus_{k,l\geq 0}\mathrm{MH}_{k,l}(G;R)$, where
$\mathrm{MH}_{k,l}(G;R)\coloneqq\mathrm{H}_{k}(\mathrm{MC}_{*,l}(G;R),\partial)$
@@
The definition of magnitude homology extends verbatim to directed graphs [Asa23].
@@
Let $G$ be a graph on the set of vertices $V(G)$. The $(k,\ell)$-eulerian magnitude chain $\mathrm{EMC}_{k,\ell}(G)$ is the free $R$-module generated by tuples $(x_{0},\dots,x_{k})\in V(G)^{k+1}$ of length $\ell$ such that $x_{i}\neq x_{j}$ for every $0\leq i,j\leq k$.
@@
Let $G$ be a (directed) graph with vertex set $V(G)$. Then, for each positive integer $\ell\geq 0$, we have
$\mathrm{EMC}_{\ast,\ell}(G)=\bigoplus_{a,b\in V(G)}\mathrm{EMC}_{\ast,\ell}(a,b)\ ,$
@@
Eulerian magnitude homology does not satisfy Künneth. In fact, let $I_{2}$ be the linear undirected graph with two vertices $x_{0},x_{1}$ and an undirected edge $\{x_{0},x_{1}\}$. Its eulerian magnitude homology is non-trivial only in bidegrees $(k,l)=(0,0),(1,1)$, in which cases it is of rank 2. The Cartesian product $I_{2}\times I_{2}$ of $I_{2}$ with itself is the cycle graph $C_{4}$ on four vertices, which has non-trivial eulerian magnitude homology in bidegree $(k,l)=(3,5)$. By dimensionality, the generators in bidegree $(3,5)$ can not come from generators of $\mathrm{EMH}_{*,*}(I_{2})\otimes\mathrm{EMH}_{*,*}(I_{2})$.
@@
A morphism of (directed) graphs $\phi\colon G_{1}\to G_{2}$ is a function $\phi\colon V(G_{1})\to V(G_{2})$ such that if $e\in E(G_{1})$ then $\phi(e)\in E(G_{2})$. Hence, a morphism of (directed) graphs sends (directed) edges to (directed) edges and, in particular, it does not allow collapsing. A morphism of (directed) graphs is called regular if it is injective (as a function). Directed graphs and regular morphisms form a category that we denote by $\mathbf{Digraph}$. Then, we have the analog of [HW17, Prop. 3.3], for family of regular morphisms of digraphs (opposed to contractions in [HW17, Prop. 3.3]):
@@
Eulerian magnitude homology is a functor on $\mathbf{Digraph}$.
@@
For a regular morphism $f\colon G^{\prime}\to G$, we let $f_{*}$ be the map defined on tuples of $G^{\prime}$ as
$f_{*}(x_{0},\dots,x_{k})=\begin{cases}(f(x_{0}),\dots,f(x_{k}))&\text{ if }\ell(f(x_{0}),\dots,f(x_{k}))=\ell(x_{0},\dots,x_{k})\\ 0&\text{ otherwise }.\end{cases}$
Since $f$ is injective on the vertices, $f_{*}$ induces a morphism on the eulerian magnitude chains, hence in eulerian magnitude homology. Furthermore, it respects the identy and compositions of regular morphisms, yielding a functor on $\mathbf{Digraph}$. ∎
We observe here that the functoriality in Proposition 1.9 can not be extended to the family of contractions. To see it, consider the undirected cyclic graph $C_{4}$ on vertices $\{x_{0},x_{1},x_{2},x_{3}\}$. Consider the contraction $c\colon C_{4}\to C_{3}$ of the edge $\{x_{2},x_{3}\}$ of $C_{4}$. As $c_{*}(x_{0},x_{1},x_{2},x_{3})=(c(x_{0}),c(x_{1}),c(x_{2}),c(x_{3}))$, the map $c_{*}$ can not be non-zero – as the tuple $(c(x_{0}),c(x_{1}),c(x_{2}),c(x_{3}))$ would have repetitions. Then, we must have $c_{*}(x_{0},x_{1},x_{2},x_{3})=0$. Likewise, we must have $c_{*}(x_{0},x_{2},x_{3})=c_{*}(x_{1},x_{2},x_{3})=0$. However, this extension to zero does not yield a chain map:
$\partial(c_{*}(x_{0},x_{1},x_{2},x_{3}))=0\neq c_{*}(\partial(x_{0},x_{1},x_{2},x_{3}))$
as $c_{*}(\partial(x_{0},x_{1},x_{2},x_{3}))=(c(x_{0}),c(x_{1}),c(x_{3}))-(c(x_{0}),c(x_{1}),c(x_{2}))$.
We provide some examples of computation of eulerian magnitude homology.
@@
Let $I_{n}$ be the linear undirected graph on $n$ vertices. Then, a direct computation shows that the maximum index $\ell$ for which $\mathrm{EMH}_{k,\ell}(I_{n})\neq 0$ is $\ell=\frac{n(n-1)}{2}$, achieved for k=n1k=n-1.
@@
Let $G=I^{n}_{2}$ be the $n$-fold Cartesian product of the graph $I_{2}$ with itself. A direct computation shows that maximum index $\ell$ for which $\mathrm{EMH}_{k,\ell}(I^{n}_{2})\neq 0$ is the maximum total Hamming distance between pairs of consecutive elements in any permutation of all $2^{n}$ binary words of length $n$ – see the A271771 sequence.
@@
[…]
and it is achieved in bidegree $(n-1,\ell(n))$. Furthermore, for every $(k,\ell)$ it holds that $\mathrm{EMH}_{k,\ell}(C_{n})$ is upper bounded by
$2\cdot\binom{n}{k+1}\binom{\ell-1}{k}\$,
@@
We consider a directed graph $G$ on the vertex set $V(G)$. For $n\geq 0$, an elementary path is an ordered sequence $(x_{0},\dots,x_{n})$ of vertices of $G$. We let $\Lambda_{n}(G)$ be the module freely generated by the $n$-elementary paths of $G$. There is a differential $\partial\colon\Lambda_{n}(G)\to\Lambda_{n-1}(G)$ as the alternating sums of the face maps forgetting the $i$-th entry of the elementary paths. An elementary path $(x_{0},\dots,x_{n})$ is called allowed if $(x_{i},x_{i+1})$ is a directed edge of $G$ for all $i=0,\dots,n-1$.
@@ 2. Regularly diagonal graphs
In this section we focus on the diagonality properties of eulerian magnitude homology of undirected graphs, providing a characterization of complete graphs in terms of eulerian magnitude homology.
@@
sending a tuple of $G$ to the tuple in $G^{\mathrm{op}}$ with same vertices and opposite order. Then, this map induces an isomorphism of magnitude homology groups [IM24, Lemma 6.3], and, in the same way, an isomorphism of eulerian magnitude homology groups. ∎
@@
If $G$ has no directed cycles, then the eulerian magnitude chain complex of $G$ is the magnitude chain complex of $G$, and the result follows. ∎
Remark 2.6.
@@
We shall see that Lemma 2.7 can be improved, and in Corollary 2.13 we provide a complete characterization of regularly diagonal undirected graphs. To proceed with such characterization, we need to analyze the relation between the existence of cycles in an undirected graph and its eulerian magnitude homology. First, we recall the following fact – see [GM24, Lemma 3.2]:
@@
Recall from Example 2.1 that undirected trees are diagonal graphs, and, as trees have no cycles, the eulerian magnitude homology in bidegree $(k,k)$, $k\geq 2$, is trivial by Proposition 2.9. One may be tempted to infer that trees are also regularly diagonal; however, this is not true, as the next example shows:
@@
Let $G$ be a finite connected regularly diagonal undirected graph, with eulerian magnitude homology groups $\mathrm{EMH}_{k,k}(G)=0$ in all bidegrees $(k,k)$ for $k>n$. Then, $G$ is the complete graph on $n+1$ vertices.
@@
  • pick any vertex $x_{0}$ of $G$ with $\ell(x_{0})=D$;
@@
Observe that, by construction, the length $l$ of the tuple $(x_{0},\dots,x_{n})$ is greater than $n$. As $(x_{0},\dots,x_{n})$ is of maximal size, it is not in the image of the differential. We want to show that it is a cycle, concluding the proof. To show that the tuple $(x_{0},\dots,x_{n})$ is a cycle, it is enough to show that all subtuples $(x_{0},\dots,\hat{x_{i}},\dots,x_{n})$ have length smaller than $(x_{0},\dots,x_{n})$. For a series of vertices $x_{i-1},x_{i},x_{i+1}$ in the tuple, observe that we always have $d_{G}(x_{i-1},x_{i+1})\leq d_{G}(x_{i-1},x_{i})+d_{G}(x_{i},x_{i+1})$ by the triangle inequality. In addition, $d_{G}(x_{i-1},x_{i+1})\leq\ell(x_{i-1})$ by definition of $\ell(x_{i-1})$, and, by construction of the tuple $(x_{0},\dots,x_{n})$, we have that $d_{G}(x_{i-1},x_{i})$ is maximal among the distances $d_{G}(x_{i-1},y)$ with $y\in V(G)\setminus\{x_{0},\dots,x_{i-1}\}$. Hence, $d_{G}(x_{i-1},x_{i})\geq d_{G}(x_{i-1},x_{i+1})$. As a consequence, we have $d_{G}(x_{i-1},x_{i+1})<d_{G}(x_{i-1},x_{i})+d_{G}(x_{i},x_{i+1})$; that is, $\mathrm{len}(x_{0},\dots,\hat{x_{i}},\dots,x_{n})<\mathrm{len}(x_{0},\dots,x_{n})$ and $(x_{0},\dots,x_{n})$ is a cycle in $\mathrm{EMH}_{n,l}(G)$. However, this is not possible because $l>n$ and $G$ is regularly diagonal. ∎
@@
If $\mathrm{gir}(G)=\gamma$ then $G$ contains $C_{\gamma}$ as a subgraph with vertex set $\{x_{1},\dots,x_{\gamma}\}$. So, it is possible to construct a tuple of vertices of type $c=(x_{1},x_{\lfloor\frac{\gamma}{2}\rfloor},x_{\gamma},x_{\lfloor\frac{\gamma}{2}\rfloor-1},\dots)$, maximizing at each step the distance between consecutive elements in $c$. Hence, by construction $c$ is a non-trivial cycle. Then, as in Example 1.14, the statement follows. ∎
@@
At this stage, having explored diagonality in $\mathrm{EMH}$ and its relationship to graph girth, a natural question arises: how does $\mathrm{EMH}(G)$ behave when $G$ is not the complete graph? Specifically, to what extent does $\mathrm{EMH}(G)$ deviate from the diagonal in such cases? To address this, we introduce a distance metric between graphs based on graph density, which quantifies how far a given graph $G$ on $n$ vertices is from the complete graph $K_{n}$. We then establish a result connecting this distance to the deviation of $\mathrm{EMH}(G)$ from the diagonal, providing a measure of how the homology of $G$ diverges from that of $K_{n}$ as the graph structure becomes sparser.
Definition 2.16.
Let $G$ be a graph. The subgraph network associated to $G$, denoted by $2^{G}$, is the undirected graph such that:
  • the set of vertices $V(2^{G})$ of $2^{G}$ is the set of all non-isomorphic connected spanning subgraphs of $G$;
  • there is an edge between two vertices of $2^{G}$, say $H_{1}$ and $H_{2}$, if and only if for every vertex $v\in V(H_{1})\cap V(H_{2})$ it holds that $|\deg_{H_{1}}(v)-\deg_{H_{2}}(v)|\leq 1$.
As the vertices of $2^{G}$ can be represented by subgraphs of $G$, we say that $H$ is contained in $2^{G}$ if it appears as a vertex of $2^{G}$. We provide in Figure 1 the subgraph network of the complete graph $K_{4}$.
0112233 2233011 3322011 2233011 3322011 2233011 2233011
Figure 1. Subgraph network of $K_{4}$. The network $2^{K_{4}}$ is represented in red, and the non-isomorphic spanning connected subgraphs of $K_{4}$ defining the vertices of $2^{K_{4}}$ are drawn in black.
The following result is a straightforward consequence of Definition 2.16.
Lemma 2.17.
Let $G$ be a connected undirected graph. Then, we have:
  1. (1)
    $2^{G}$ is connected;
  2. (2)
    the diameter of $2^{G}$ is $\max_{v\in V(G)}\deg_{G}(v)$.
Let $G$ and $H$ be undirected graphs with $n$ vertices. We further define the covering subgraph network of $G$ and $H$, denoted by $2^{(G,H)}$, as the subgraph network associated to the complete graph $K_{n}$.
Example 2.18.
Let $G$ be the linear graph $I_{4}$ on $4$ vertices, and let $H$ be the bipartite graph $K_{1,3}$. Then, the covering subgraph network $2^{(I_{4},K_{1,3})}$ is the subgraph network $2^{K_{4}}$ represented in Figure 1.
We are now ready to introduce a metric between graphs with same number of vertices. To do it, recall from Lemma 2.17 that the graph $2^{(G,H)}$ is connected.
Definition 2.19.
Let $G,H$ be undirected graphs with $|V(G)|=|V(H)|$. The distance $\Delta$ between $G$ and $H$, is the distance $\Delta(G,H)$ between $G$ and $H$ considered as vertices of $2^{(G,H)}$:
$\Delta(G,H)\coloneqq d_{2^{(G,H)}}(G,H)\ ,$
where $d_{2^{(G,H)}}$ is the path metric on $2^{(G,H)}$.
Example 2.20.
The distance $\Delta(K_{1,3},C_{4})$ between the bipartite graph $K_{1,3}$ and the cycle graph $C_{4}$ is $\Delta(K_{1,3},C_{4})=1$, as shown in Figure 1.
Notation 2.21.
For the undirected simple complete graph $K_{n}$, we call $\gamma_{n}^{s}$ the girth of the maximal sub-cycle achievable by removing $s$ edges from $K_{n}$, see Figure 2 for an example.
0123456789
Figure 2. By removing $34$ edges from $K_{10}$ we can build a $C_{6}$, so we have $\gamma_{10}^{34}=6$.
With these preliminaries in place, we are now ready to state the result connecting $\Delta(G,K_{n})$ to the deviation of $\mathrm{EMH}(G)$ from the diagonal.
Proposition 2.22.
Let $G$ be a graph on $n$ vertices such that $\Delta(G,K_{n})=h$ and $C_{\gamma}$, with $\gamma\geq\gamma^{h}_{n}$, is a cyclic subgraph of $G$. Then $G$ has non-trivial eulerian magnitude homology until at least subdiagonal $\lceil\frac{(\gamma-1)^{2}}{2}\rceil-(\gamma-1)$.
Proof.
The proof is a direct application of Lemma 2.14. ∎
The hypothesis in Proposition 2.22 that $G$ contains a big enough cycle $C_{\gamma}$ is crucial: it is not enough for the graph $G$ to be globally far from the complete graph in terms of edge count – the distribution of missing edges plays a critical role:
Example 2.23.
Consider the graph $G=K_{n-1}\cup\{e\}$, where $e$ is a single edge connecting one isolated vertex $v\notin K_{n-1}$ to a vertex in $K_{n-1}$. Then $\Delta(G,K_{n})=n-2$, since it lacks all but one of the n1n-1 edges incident to $v$. Nevertheless, a direct computation shows that $\mathrm{EMH}(G)$ is non-trivial only until the first subdiagonal.
This example illustrates that a large edge deficit is not sufficient, on its own, to force non-trivial $\mathrm{EMH}(G)$ strictly below the first subdiagonal: the missing edges must induce a “structurally sparse” region in the graph, such as a large cycle.